OpenStudy (anonymous):

find all solutions of the equation sinx+2sin^2 x =0

OpenStudy (anonymous):

let \[sinx=y\]

OpenStudy (anonymous):

\[2y ^{2}+y=0\]

OpenStudy (anonymous):

\[Y(2y+1)=0\]

OpenStudy (anonymous):

\[y=0\]

OpenStudy (anonymous):

\[2y+1=0\]

OpenStudy (anonymous):

\[y=\frac{ -1 }{ 2}\]

OpenStudy (anonymous):

will that be on quad 2?

OpenStudy (anonymous):

\[sinx=0 and sinx=\frac{ -1 }{ 2 }\]

OpenStudy (anonymous):

then you find your values

OpenStudy (anonymous):

how do you find the value of sinx? please help me out.. this will help me for upcoming test

OpenStudy (anonymous):

\[\sin ^{-1}(0)=0\]

OpenStudy (anonymous):

sinx=-1/2 is pi/6 right ?

OpenStudy (anonymous):

@Directrix please can you help him

OpenStudy (anonymous):

any help with finding the value of sinx=-1/2

Directrix (directrix):

sinx+2sin^2 x =0 2 sin² x + sinx = 0 sinx (2 sinx + 1) = 0 sin x = 0 OR 2 sin x + 1 = 0 sin x = 0 Or 2 sin x = -1 sin x = 0 or sin x = -1/2

Directrix (directrix):

What is the solution domain? Zero to 360 degrees or 0 to 2 pi radians or all Real numbers?

OpenStudy (anonymous):

it says all solutions, so im assuming all real numbers

Directrix (directrix):

You will need to learn the fundamental trig values.

OpenStudy (anonymous):

so it will be -30 right?

Directrix (directrix):

sin x = 0 x = ..., -2π, -π, 0, π, 2π,...

Directrix (directrix):

or sin x = -1/2 ..... -5π/6 ,-π/6, 7π/6, 11π/6, ...

Directrix (directrix):

Let's think how to write these in a less messy way: x = ..., -2π, -π, 0, π, 2π,... Union x = ..... -5π/6 ,-π/6, 7π/6, 11π/6, ...

OpenStudy (anonymous):

by messy, you mean writing it in degrees?? ...... sorry for asking to much... im just not good at this chapter and i really need help

Directrix (directrix):

See attached file. Z is the set of integers. http://www.wolframalpha.com/input/?i=sinx%2B2sin%5E2+x+%3D0

Directrix (directrix):

By messy, I mean to write then not individually but in terms of a multiple of an integer. There are are infinitely many solutions.

Directrix (directrix):

The first thing that went wrong is that you divided both sides of an equation by a variable and lost half the answers. See attached file.

Directrix (directrix):

The next problem is leaning the values of the trig functions at 0, pi/6, pi/4, pi/3, and pi/2,

OpenStudy (anonymous):

i factored it , but is there a better way to do it without losing solutions ?

OpenStudy (anonymous):

so what other options do i have besides factoring ?

Directrix (directrix):

If the expression will factor, do that. You can take y² = y and divide both sides by y to get y = 1 IF you also remember that y could be equal to zero. I always factor. If such an equation is quadratic in degree and will not factor, you can always use the Quadratic Formula.

OpenStudy (anonymous):

so what is the proper way to solve this equation? im confused now.. is telijahmed's work wrong? if so, how do you solve it ?

Directrix (directrix):

This is correct work.

Directrix (directrix):

Solutions

OpenStudy (anonymous):

ok, and from that how do you find all the value of sinx=-1/2

OpenStudy (anonymous):

please help.....